# Detecting dark objects with plasma microlensing by their gravitational wakes

### v1

@@ 1 Introduction
Plasma microlensing can gravitationally probe the same objects if they reside in an (at least partially) ionized medium. For gravitational lensing in the weak-field regime, the perturbation $\phi(\bm{x})\equiv 1-n_{\mathrm{r}}(\bm{x})$ to the refractive index $n_{\mathrm{r}}(\bm{x})$ is
$\phi_{G}(\bm{x})=\frac{2}{c^{2}}\Phi(\bm{x}),$ (gravitational lensing) (1)
where $\Phi$ is the Newtonian gravitational potential and $\bm{x}$ is position. Meanwhile, for plasma lensing of light of wavelength $\lambda$, the refractive index is perturbed by
$\phi_{\mathrm{p}}(\bm{x})=\frac{1}{2\pi}r_{e}\lambda^{2}n_{e}(\bm{x})$ (plasma lensing) (2)
(e.g. Wagner & Er, 2020), where $r_{e}\simeq 2.818\times 10^{-13}~{}\mathrm{cm}$ is the classical electron radius and $n_{e}$ is the number density of electrons. But a gravitational potential $\Phi$ induces variations $\delta n_{e}(\bm{x})=n_{e}(\bm{x})-\bar{n}_{e}$ that are of order
$\frac{\delta n_{e}(\bm{x})}{\bar{n}_{e}}\sim-\frac{\Phi(\bm{x})}{c_{\mathrm{s}}^{2}},$ (3)
where $c_{\mathrm{s}}$ is the sound speed of the plasma and $\bar{n}_{e}$ is the mean electron density. [^1] For a stationary object with $|\Phi|\ll c_{\mathrm{s}}^{2}$, equation (3) is exact. For the same gravitational source object, the ratio between the effects of plasma and gravitational lensing is therefore of order
$\frac{\phi_{\mathrm{p}}}{\phi_{G}}$ $\sim-\frac{r_{e}\lambda^{2}\bar{n}_{e}}{4\pi}\frac{c^{2}}{c_{\mathrm{s}}^{2}}$
@@
This comparison suggests that plasma lensing is competitive with gravitational lensing only at very low radio frequencies. However, for a moving object, the spatial structure of the plasma lens is very different from that of the gravitational lens, and so the signals from plasma and gravitational microlensing can be very different. This work quantifies the plasma lens associated with a gravitational wake and characterizes the plasma microlensing signatures that it can produce.
@@ 3.2 Wakes as one-dimensional lenses
Figure 5: Structure of the edge of the gravitational wake of a dark object of mass $30~{}\mathrm{M}_{\odot}$ moving at $200~{}\mathrm{km}\,\mathrm{s}^{-1}$. For a range of distances $l$ from the object (different colors), we show the excess column density $N_{e}$ as a function of the perpendicular distance $l_{\perp}$ from the edge of the Mach cone associated with the wake. $l_{\perp}$ is defined such that $l_{\perp}<0$ lies outside the cone, and we show it in units of the acoustic damping scale $\sigma_{\mathrm{coll}}$ (which depends on $l$). We show both the collisionless and the damped acoustic descriptions. For comparison, the dotted line indicates the simplified description of equation (27), which closely matches the damped acoustic approximation at large $l$.
@@ 3.5 Known plasma microlensing events
Plasma microlensing events have been observed (e.g. Fiedler et al., 1987, 1994; Coles et al., 2015), and the attribution of these “extreme scattering events” remains unclear (e.g. Clegg et al., 1998; Dong et al., 2018; Shi & Xu, 2021; Jow et al., 2024). It is appropriate to discuss whether these events could be attributed to the gravitational wakes of dark objects. Models that reproduce these events tend to involve an electron column density $N_{e}$ that rises and then falls again (Clegg et al., 1998; Dong et al., 2018; Jow et al., 2024). This is different from the picture of a gravitational wake developed in figure 5, in which $N_{e}$ rises but does not fall. Also, the few-month time scales associated with extreme scattering events are much shorter than those associated with gravitational wakes in the far regime that we have focused on.

### v2

@@ 1 Introduction
Plasma microlensing can gravitationally probe the same objects if they reside in an (at least partially) ionized medium. In plasma lensing, radio-frequency light is deflected by gradients in the density of free electrons (e.g. Romani et al., 1987; Clegg et al., 1998; Tuntsov et al., 2016; Dong et al., 2018; Shi & Xu, 2021). But the gravitational influence of an invisible massive object would induce density variations in the plasma, giving rise to plasma lensing as a purely gravitational signature of the object’s presence.
We can make a simple estimate of the relative impacts of gravitational and plasma lensing arising from the gravitational potential $\Phi(\bm{x})$ of the same massive object. Let the refractive index be $n_{\mathrm{r}}(\bm{x})=1-\phi^{(\text{g})}(\bm{x})-\phi^{(\text{p})}(\bm{x})$, where $\phi^{(\text{g})}$ and $\phi^{(\text{p})}$ are small gravitational and plasma contributions, respectively. These contributions are
$\phi^{(\text{g})}(\bm{x})$ $=\frac{2}{c^{2}}\Phi(\bm{x})$ (gravitational lensing) (1)
$\phi^{(\text{p})}(\bm{x})$ $=\frac{1}{2\pi}r_{e}\lambda^{2}n_{e}(\bm{x})$ (plasma lensing) (2)
(e.g. Wagner & Er, 2020), where $\lambda$ is the light wavelength, $r_{e}\simeq 2.818\times 10^{-13}~{}\mathrm{cm}$ is the classical electron radius, and $n_{e}$ is the number density of electrons. But the potential $\Phi$ induces plasma density variations $\delta(\bm{x})=[n_{e}(\bm{x})-\bar{n}_{e}]/\bar{n}_{e}$ that are of order
$\delta\sim-\Phi/c_{\mathrm{s}}^{2},$ (3)
where $c_{\mathrm{s}}$ is the sound speed of the plasma and $\bar{n}_{e}$ is the mean electron density.111 For a stationary object with $|\Phi|\ll c_{\mathrm{s}}^{2}$, equation (3) is exact. This can be seen by comparing the wave equation for $\delta$ driven by a mass density $\rho$, $-\partial^{2}\delta/\partial t^{2}+c_{\mathrm{s}}^{2}\nabla^{2}\delta=-4\pi G\rho$, with the Poisson equation for $\Phi$, $\nabla^{2}\Phi=4\pi G\rho$. If all time dependence is suppressed, evidently $\delta=-\Phi/c_{\mathrm{s}}^{2}$ up to additive shifts. The ratio between the effects of plasma and gravitational lensing is therefore of order
$\frac{\phi^{(\text{p})}}{\phi^{(\text{g})}}$ $\sim-\frac{r_{e}\lambda^{2}\bar{n}_{e}}{4\pi}\frac{c^{2}}{c_{\mathrm{s}}^{2}}$
@@
This comparison suggests that plasma lensing is competitive with gravitational lensing only at very low radio frequencies. However, an object moving supersonically induces plasma density variations $\delta$ in the form of an extended gravitational wake (Ostriker, 1999), whose spatial shape can be very different from that of the potential $\Phi$ (c.f. equation 3). That is, the spatial structure of the plasma lens can be very different from that of the gravitational lens, and so the signals from plasma and gravitational microlensing can be very different. This work quantifies the plasma lens associated with a gravitational wake and characterizes the plasma microlensing signatures that it can produce.
@@ 3.2 Wakes as one-dimensional lenses
Figure 5: Structure of the edge of the gravitational wake of a dark object of mass $30~{}\mathrm{M}_{\odot}$ moving at $200~{}\mathrm{km}\,\mathrm{s}^{-1}$. For a range of distances $l$ from the object (different colors), we show the excess column density $N_{e}$ as a function of the perpendicular distance $l_{\perp}$ from the edge of the Mach cone associated with the wake. $l_{\perp}$ is defined such that $l_{\perp}<0$ lies outside the cone, and we show it in units of the acoustic damping scale $\sigma_{\mathrm{coll}}$ (which depends on $l$). We show both the damped acoustic description (solid curves) and the collisionless description (dashed curves, only shown for $l\leq 1000~{}\mathrm{au}$). For comparison, the dotted line indicates the simplified description of equation (27), which closely matches the damped acoustic approximation at large $l$.
@@ 3.5 Known plasma microlensing events
Plasma microlensing events have been observed (e.g. Fiedler et al., 1987, 1994; Coles et al., 2015; Bannister et al., 2016), and the attribution of these “extreme scattering events” remains unclear (e.g. Clegg et al., 1998; Dong et al., 2018; Shi & Xu, 2021; Jow et al., 2024). It is appropriate to discuss whether these events could be attributed to the gravitational wakes of dark objects. Models that reproduce these events tend to involve an electron column density $N_{e}$ that rises and then falls again (Clegg et al., 1998; Dong et al., 2018; Jow et al., 2024). This is different from the picture of a gravitational wake developed in figure 5, in which $N_{e}$ rises but does not fall. Also, the few-month time scales associated with extreme scattering events are much shorter than those associated with gravitational wakes in the far regime that we have focused on.